OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at y = ±5/4x.

OpenStudy (anonymous):

asymptotes \[y=\pm \frac{b}{a}x\]

OpenStudy (anonymous):

sorry that should be a - sign on top

OpenStudy (anonymous):

x^2/16 - y^2/25 =1 ?

OpenStudy (anonymous):

\[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

what do u think?

OpenStudy (anonymous):

I think it's right

OpenStudy (anonymous):

i am not sure heehehe

OpenStudy (anonymous):

:( come on now

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

so it's right?

OpenStudy (anonymous):

u win

OpenStudy (anonymous):

YAY thank you!

OpenStudy (anonymous):

ur welcome!